OpenStudy (alli6106):

There are 60 students in a speech contest. Yesterday, 2/3 of them gave their speeches. Today, 4/5 of the remaining students gave their speeches. How many students still haven't given their speeches?

OpenStudy (triciaal):



OpenStudy (anonymous):

Yesterday, 40 students gave their speeches: \[\frac{ 2 }{ 3 }\times60=40\] Which left 20 students who hadn't given their speeches: \[60-40=20\] Today, 16 students gave their speeches: \[\frac{ 4 }{ 5 }\times20=16\] Which leaves 4 students who still haven't given their speeches: \[20-16=4\]